Is the ordered pair (3,-4) a solution to: 5x-2y=17?
Is the ordered pair (-2,8) a solution to: y-9=3x?
Is the ordered pair (-1,3) a solution to: -3x+4y=15?
Is the ordered pair (1,-2) a solution to: -5x-3y=1?
Is the ordered pair (-2,7) a solution to: 2y+6x=2?
Is the ordered pair (4,6) a solution to: 2y+6x=2?
Is the ordered pair (9,-3) a solution to: 4y-3x=45?
Is the ordered pair (-7,2) a solution to: 19+2x=2y?

HELP!!! i wasnt here the day we learned this in class and now i have homework and i need to show work so PLEASE HURRY!!! i am doing a year up course in math so it is complicated (for me)and i am desperate for help! SHOW WORK PLEASE!!!

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

(1)(3,-4) 5x-2y=17========NO
5x+-2y+2y=17+2y
5x+0=17+2y
5x=17+2y
x=3.4+0.4y
(2)no (3) no (4)yes (5)no
(6)no (7)no (8) no
OK, NOW I ONLY WORKED OUT THE FIRST PROBLEM
YOU SHOULD BE ABLE TO LOOK AT THIS EXAMPLE
AND SEE HOW TO UNDERSTAND HOW TO DO
THE OTHER PROBLEMS

Answer by Alan3354(69443)

Is the ordered pair (3,-4) a solution to: 5x-2y=17?
Sub 3 for x, and -4 for y and see if it's equal.
5x-2y=17
5*3 - 2*(-4) =? 17
15 + 8 =? 17
23 =? 17
Not equal, so not a solution.
-------------------------------
All the others are the same. Try them. Email via the Thank You note for help, or to check your work.
===============================================
Is the ordered pair (-2,8) a solution to: y-9=3x?
Is the ordered pair (-1,3) a solution to: -3x+4y=15?
Is the ordered pair (1,-2) a solution to: -5x-3y=1?
Is the ordered pair (-2,7) a solution to: 2y+6x=2?
Is the ordered pair (4,6) a solution to: 2y+6x=2?
Is the ordered pair (9,-3) a solution to: 4y-3x=45?
Is the ordered pair (-7,2) a solution to: 19+2x=2y?